A firm has estimated the two-month cash budget below. What is the cash surplus or deficit for these two months? (Enter your answers in millions of dollars.)

($ in millions) March April
Sales 93.0 86.0
Cash collection 81.0 83.0
Total cash disbursement 69.0 66.0
Net cash flow 12.0 17.0
Cumulative net cash flow 29.0
Minimum cash balance 20.0 20.0
Cash surplus or deficit
Answer #1

April's Cumulative net cash flow = March's Cumulative net cash flow + Net Cash Flow

= 29 + 17 = $46 million

March Cash Surplus or Deficit = Cumulative net cash flow - Minimum cash balance

= 29 - 20 = $9 million

April Cash Surplus or Deficit = Cumulative net cash flow - Minimum cash balance

= 46 - 20 = $26 million